# Two-Stage Nuisance Function Estimation for Causal Mediation Analysis

### v2

@@ 4 Proposed Estimation Approach
We define the following nuisance functions.
$\pi_{1}(X)=\frac{1}{p(A=0\mid X)},$ $\forall X\in\mathcal{X},$
$\pi_{2}(M,X)=\frac{p(M\mid A=0,X)}{p(M,A=1\mid X)},$ $\forall M\in\mathcal{M},X\in\mathcal{X},$
$\mu_{1}(M,X)=\mathbb{E}[Y\mid A=1,M,X],$ $\forall M\in\mathcal{M},X\in\mathcal{X},$
$\mu_{2}(X)=\mathbb{E}[\mu_{1}(M,X)\mid A=0,X],$ $\forall X\in\mathcal{X}.$
@@
$\psi_{0}=\mathbb{E}[\mu_{2}(X)].$
We start by presenting the steps of our proposed estimation method, which contains cross-fitting (Chernozhukov et al.,, 2018) for separating the samples used for estimating the nuisance functions from that used for estimating the parameter of interest.
Estimating the nuisance functions:
We partition the samples into $L$ equal size folds $\{I_{1},\dots,I_{L}\}$. For $\ell\in\{1,\dots,L\}$, we estimate the nuisance functions $(\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})$ on data from all folds but $I_{\ell}$, which we denote as $I_{\ell}^{c}$. In order for estimating the nuisance functions, we proceed in the following two stages.
@@
      $\mathbb{E}\left[\left\{(1-A)\hat{\pi}_{1,\ell}(X)-1\right\}h_{1}(X)\right]\approx 0,~{}~{}~{}~{}~{}~{}\forall h_{1}\in\mathcal{H}_{1},$ (1)
@@
      $\mathbb{E}\left[\left\{A\hat{\pi}_{2,\ell}(M,X)-(1-A)\hat{\pi}_{1,\ell}(X)\right\}h_{2}(M,X)\right]\approx 0,~{}~{}~{}~{}~{}~{}\forall h_{2}\in\mathcal{H}_{2},$ (2)
@@
    • Given $\hat{\mu}_{1,\ell}$ from Stage 1, we estimate $\mu_{2}$ by regressing $\hat{\mu}_{1,\ell}$ on $\{X,A=0\}$ to obtain $\hat{\mu}_{2,\ell}$.
Estimating the parameter of interest:
For any estimates of nuisance functions $\{\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}\}$, let
@@
For all $\ell$, let $\hat{\psi}^{mr}_{\ell}$ be the estimator of $\psi_{0}$ defined as,
$\hat{\psi}^{mr}_{\ell}=\mathbb{E}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right],$
@@
We propose a non-parametric estimator for Stage 1 based on minimax optimization. A naive approach to implement this would be by solving the following optimization problem.
$\tilde{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right],$
@@
$\breve{\pi}_{1}=\arg\min_{\pi_{1}\in L^{2}(P_{0,X})}\max_{h_{1}\in L^{2}(P_{0,X})}\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right].$
Inspired by Proposition 2, we estimate $\pi_{1}$ as
$\hat{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]+R_{n}^{1}(\pi_{1},h_{1}),$ (3)
@@
Our proposed nonparametric estimator $\hat{\pi}_{1}$ in Equation (3) can also be used in the setting of Imai and Ratkovic, (2014) for estimating average treatment effect. Therefore, this provides a nonparametric version of the their covariate balancing propensity score (CBPS) method.
We can have similar arguments regarding $\pi_{2}$ in Stage 2.
@@
$\breve{\pi}_{2}=\arg\min_{\pi_{2}\in L^{2}(P_{0,(M,X)})}\max_{h_{2}\in L^{2}(P_{0,(M,X)})}\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right].$
@@
$\hat{\pi}_{2}=\arg\min_{\pi_{2}\in\Pi_{2}}\max_{h_{2}\in\mathcal{H}_{2}}\mathbb{E}_{n}\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]+R_{n}^{2}(\pi_{2},h_{2}),$ (4)
@@
In order to estimate $\mu_{1}$ in Stage 1 and $\mu_{2}$ in Stage 2 of the approach, any standard parametric or non-parametric regression method can be used.
4.1 Doubly Robust Kernel Estimator
In this section, we describe solving the minimax problems (3) and (4) when RKHSes are used as the hypothesis classes. We use the method adopted by Ghassami et al., (2022), which builds upon the non-parametric estimation approach proposed by Dikkala et al., (2020) based on adversarial learning for solving integral equations within the semi-parametric proximal causal inference framework.
Let $\Pi_{1}$, $\mathcal{H}_{1}$, $\Pi_{2}$, and $\mathcal{H}_{2}$ be RKHSes with kernels $K_{\Pi_{1}}$, $K_{\mathcal{H}_{1}}$, $K_{\Pi_{2}}$ and $K_{\mathcal{H}_{2}}$, respectively, equipped with the RKHS norms $\|\cdot\|_{\Pi_{1}}$, $\|\cdot\|_{\mathcal{H}_{1}}$, $\|\cdot\|_{\Pi_{2}}$ and $\|\cdot\|_{\mathcal{H}_{2}}$. We propose the following Tikhonov regularization-based optimizations:
$\hat{\pi}_{1}$ $=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]-\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}+\lambda_{\Pi_{1}}\|\pi_{1}\|_{\Pi_{1}}^{2},$ (5)
$\hat{\pi}_{2}$ $=\arg\min_{\pi_{2}\in\Pi_{2}}\max_{h_{2}\in\mathcal{H}_{2}}\mathbb{E}_{n}\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
@@ 4.1 Doubly Robust Kernel Estimator
where $\lambda_{\mathcal{H}_{1}}$, $\lambda_{\Pi_{1}}$, $\lambda_{\mathcal{H}_{2}}$, and $\lambda_{\Pi_{2}}$ are regularization hyperparameters that are decreasing in $n$.
For the convergence analysis of the proposed minimax estimators, we direct the reader to (Dikkala et al.,, 2020; Ghassami et al.,, 2022).
@@
$\alpha=-\Big{(}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}+n^{2}\lambda_{\Pi_{1}}K_{\Pi_{1},n}\Big{)}^{\dagger}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}e_{n},$
where $\Gamma_{1}=K_{\mathcal{H}_{1},n}(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}$, $diag(1-A)$ is a diagonal matrix with $1-A_{i}$ as the $i$-th diagonal entry, and $e_{n}:=\begin{pmatrix}1&1&\cdots&1\end{pmatrix}^{\top}$, and $\dagger$ denotes Moore-Penrose pseudo-inverse.
@@
$\beta=-\Big{(}K_{\Pi_{2},n}diag(A)\Gamma_{2}diag(A)K_{\Pi_{2},n}+n^{2}\lambda_{\Pi_{2}}K_{\Pi_{2},n}\Big{)}^{\dagger}K_{\Pi_{2},n}diag(A)\Gamma_{2}\big{(}(1-A)\hat{\pi}_{1}(X)\big{)}_{n},$
where $\Gamma_{2}=K_{\mathcal{H}_{2},n}(\frac{1}{n}K_{\mathcal{H}_{2},n}+\lambda_{\mathcal{H}_{2}}I_{n})^{-1}$, $diag(A)$ is a diagonal matrix with $A_{i}$ as the $i$-th diagonal entry, and $\big{(}(1-A)\hat{\pi}_{1}(X)\big{)}_{n}:=\begin{pmatrix}(1-A_{1})\hat{\pi}_{1}(X_{1})&(1-A_{2})\hat{\pi}_{1}(X_{2})&\cdots&(1-A_{n})\hat{\pi}_{1}(X_{n})\end{pmatrix}^{\top}$.
@@ 5 Theoretical Analysis
We start by making the observation that the true nuisance functions $\pi_{1}$ and $\pi_{2}$ satisfy conditions (1) and (2) with equality.
@@
    $\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]=0,~{}~{}~{}~{}~{}~{}\forall h_{1}\in L^{2}(P_{0,X}).$
@@
    $\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]=0,~{}~{}~{}~{}~{}~{}\forall h_{2}\in L^{2}(P_{0,(M,X)}).$
Suppose conditions (1) and (2) hold with equality and $\mathcal{H}_{1}=L^{2}(P_{0,X})$ and $\mathcal{H}_{2}=L^{2}(P_{0,(M,X)})$. Then, it is easy to see that $\hat{\pi}_{1}=\pi_{1}$ and $\hat{\pi}_{2}=\pi_{2}$ with probability one. However, when $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are not the corresponding $L^{2}$ spaces, we can have conditions (1) and (2) hold with equality even though $\hat{\pi}_{1}\neq\pi_{1}$ and $\hat{\pi}_{2}\neq\pi_{2}$. Nevertheless, satisfying conditions (1) and (2) can be sufficient for controlling the bias of the final estimator of the parameter of interest. We formalize this argument by looking at the structure of the bias of $\hat{\psi}^{mr}$.
@@
The bias of the estimator $\hat{\psi}^{mr}$ satisfies the following equality:
$\mathbb{E}[\hat{\psi}^{mr}]-\psi_{0}$ $=\mathbb{E}\left[\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
From the structure of the bias of the estimator $\hat{\psi}^{mr}$, it is evident that if there exists $h_{1}\in\mathcal{H}_{1}$ which well-approximates $\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}$, and there exists $h_{2}\in\mathcal{H}_{2}$ which well-approximates $\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}$, and further conditions (1) and (2) hold, then we can eliminate the bias of the estimator $\hat{\psi}^{mr}$ even if we do not have a perfect fit of $\pi_{1}$ and $\pi_{2}$. This is in fact the main motivation of our two-stage approach. It substitutes fitting $p(A=1\mid X)$ and $p(M\mid A,X)$, which can be a difficult task, by a much weaker requirement based on balancing.
Importantly, under the following positivity assumption, our approach retains the multiple-robustness property, as stated in Proposition 6.
@@
For some $\epsilon>0$, we have $p(A=0\mid X)>\epsilon$ and $p(A=1\mid M,X)>\epsilon$.
@@
Under Assumption 2 , $\hat{\psi}^{mr}$ is asymptotically unbiased if at least one of the following conditions holds:.
@@
Figure 2 summarizes the result in Proposition 6. Note that our estimator $\hat{\psi}^{mr}$ is based on an influence function, which always leads to a second-order bias.
$\hat{\pi}_{1}$$\hat{\pi}_{2}$$\hat{\mu}_{1}$$\hat{\mu}_{2}$tage 1:Stage 2:

### v3

@@ 4 Proposed Estimation Approach
4.1 Two-Stage Nuisance Function Estimation
In this section, we present our two-stage nuisance function estimation approach. We start by introducing a reparametrization of the likelihood function based on the following nuisance functions.
$\pi_{1}(X=\cdot):=\frac{1}{p(A=0\mid X=\cdot)},$
$\pi_{2}(M=\cdot,X=\cdot):=\frac{p(M=\cdot\mid A=0,X=\cdot)}{p(M=\cdot,A=1\mid X=\cdot)},$
$\mu_{1}(M=\cdot,X=\cdot):=E[Y\mid A=1,M=\cdot,X=\cdot],$
$\mu_{2}(X=\cdot):=E[\mu_{1}(M,X)\mid A=0,X=\cdot].$
@@ 4.1 Two-Stage Nuisance Function Estimation
$\psi_{0}=E[\mu_{2}(X)].$
We start by presenting the steps of our proposed estimation method, which contain cross-fitting (Chernozhukov et al.,, 2018) for separating the samples used for estimating the nuisance functions from that used for estimating the parameter of interest.
Estimating the nuisance functions. We partition the samples into $L$ folds $\{I_{1},\dots,I_{L}\}$ with roughly equal size $m=n/L$. For $\ell\in\{1,\dots,L\}$, we estimate the nuisance functions as $(\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})$ on data from all folds but $I_{\ell}$, which we denote as $I_{\ell}^{c}$. In order to estimate the nuisance functions, we proceed in the following two stages.
@@
    $E\left[\left\{(1-A)\hat{\pi}_{1,\ell}(X)-1\right\}h_{1}(X)\right]\approx 0,~~~~~~\forall h_{1}\in\mathcal{H}_{1},$ (1)
@@
    $E\left[\left\{A\hat{\pi}_{2,\ell}(M,X)-(1-A)\hat{\pi}_{1,\ell}(X)\right\}h_{2}(M,X)\right]\approx 0,~~~~~~\forall h_{2}\in\mathcal{H}_{2},$ (2)
@@
  2. (ii)
    Given $\hat{\mu}_{1,\ell}$ from Stage 1, we estimate $\mu_{2}$ by regressing $\hat{\mu}_{1,\ell}(X,M)$ on $\{X,A=0\}$ to obtain $\hat{\mu}_{2,\ell}$.
Estimating the parameter of interest. For any choice of nuisance functions $\{\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}\}$, let
@@
For all $\ell$, let $\hat{\psi}^{\text{2S}}_{\ell}$ be the estimator of $\psi_{0}$ defined as,
$\hat{\psi}^{\text{2S}}_{\ell}=E_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right],$
@@
We propose a nonparametric estimator for Stage 1 based on minimax optimization. To make the notation simple, throughout the rest of Section 4, assume we use a sample of size $n$ for estimating the nuisance functions. A direct implementation is to solve the following optimization problem.
$\tilde{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}E_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right],$
@@
$\breve{\pi}_{1}=\arg\min_{\pi_{1}\in L^{2}(P_{0,X})}\max_{h_{1}\in L^{2}(P_{0,X})}E\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right].$
Inspired by Proposition 2, we estimate $\pi_{1}$ as,
$\hat{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}E_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]+R_{n}^{1}(\pi_{1},h_{1}),$ (3)
@@
Imai and Ratkovic, (2014) proposed the covariate balancing propensity score (CBPS) method for estimating the average treatment effect. Our proposed technique for estimating $\hat{\pi}_{1}$ in Equation (3) can be viewed as a nonparametric version of their covariate balancing approach. Therefore, our proposed nonparametric estimator $\hat{\pi}_{1}$ can also be used in the setting of Imai and Ratkovic, (2014) for estimating average treatment effect without resorting to parametric models for the inverse of the propensity score.
We have similar arguments regarding $\pi_{2}$ in Stage 2, which lead to the following result.
@@
$\breve{\pi}_{2}=\arg\min_{\pi_{2}\in L^{2}(P_{0,(M,X)})}\max_{h_{2}\in L^{2}(P_{0,(M,X)})}E\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right].$
@@
$\hat{\pi}_{2}=\arg\min_{\pi_{2}\in\Pi_{2}}\max_{h_{2}\in\mathcal{H}_{2}}E_{n}\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]+R_{n}^{2}(\pi_{2},h_{2}),$ (4)
@@
In order to estimate $\mu_{1}$ in Stage 1 and $\mu_{2}$ in Stage 2 of the approach, one may use any standard parametric or nonparametric regression method.
4.2 A Kernel Method-Based Implementation
In this section, we describe solving the minimax problems (3) and (4) when RKHSes are used as the hypothesis classes. We use the method of Ghassami et al., (2022), which extends the nonparametric estimation approach proposed by Dikkala et al., (2020) for solving integral equations via adversarial learning within the semiparametric proximal causal inference framework.
Let $\Pi_{1}$, $\mathcal{H}_{1}$, $\Pi_{2}$, and $\mathcal{H}_{2}$ be RKHSes with kernels $K_{\Pi_{1}}$, $K_{\mathcal{H}_{1}}$, $K_{\Pi_{2}}$, and $K_{\mathcal{H}_{2}}$, respectively, equipped with the RKHS norms $\|\cdot\|_{\Pi_{1}}$, $\|\cdot\|_{\mathcal{H}_{1}}$, $\|\cdot\|_{\Pi_{2}}$, and $\|\cdot\|_{\mathcal{H}_{2}}$. We propose the following Tikhonov regularization-based optimization problems:
$\hat{\pi}_{1}$ $=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}E_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]-\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}+\lambda_{\Pi_{1}}\|\pi_{1}\|_{\Pi_{1}}^{2},$ (5)
$\hat{\pi}_{2}$ $=\arg\min_{\pi_{2}\in\Pi_{2}}\max_{h_{2}\in\mathcal{H}_{2}}E_{n}\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
@@ 4.2 A Kernel Method-Based Implementation
where $\lambda_{\mathcal{H}_{1}}$, $\lambda_{\Pi_{1}}$, $\lambda_{\mathcal{H}_{2}}$, and $\lambda_{\Pi_{2}}$ are regularization hyperparameters that are decreasing in $n$. The convergence analysis of the proposed minimax estimators can be done similar to the analysis presented in (Dikkala et al.,, 2020; Ghassami et al.,, 2022), and hence omitted here.
@@
    $\alpha=\Big(K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}+n^{2}\lambda_{\Pi_{1}}K_{\Pi_{1},n}\Big)^{\dagger}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}e_{n},$
    where $\Gamma_{1}=\frac{1}{4}K_{\mathcal{H}_{1},n}(\frac{1}{4n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}$, diag(1A)diag(1-A) is a diagonal matrix with $1-A_{i}$ as the $i$-th diagonal entry, $e_{n}:=\begin{pmatrix}1&1&\cdots&1\end{pmatrix}^{\top}$, and $\dagger$ denotes Moore-Penrose pseudo-inverse.
  2. (b)
@@
    $\beta=\Big(K_{\Pi_{2},n}diag(A)\Gamma_{2}diag(A)K_{\Pi_{2},n}+n^{2}\lambda_{\Pi_{2}}K_{\Pi_{2},n}\Big)^{\dagger}K_{\Pi_{2},n}diag(A)\Gamma_{2}\big((1-A)\hat{\pi}_{1}(X)\big)_{n},$
    where $\Gamma_{2}=\frac{1}{4}K_{\mathcal{H}_{2},n}(\frac{1}{4n}K_{\mathcal{H}_{2},n}+\lambda_{\mathcal{H}_{2}}I_{n})^{-1}$, $diag(A)$ is a diagonal matrix with $A_{i}$ as the $i$-th diagonal entry, and $\big((1-A)\hat{\pi}_{1}(X)\big)_{n}:=\begin{pmatrix}(1-A_{1})\hat{\pi}_{1}(X_{1})&(1-A_{2})\hat{\pi}_{1}(X_{2})&\cdots&(1-A_{n})\hat{\pi}_{1}(X_{n})\end{pmatrix}^{\top}$.
@@ 5 Theoretical Analysis
5.1 Bias Analysis and Robustness Properties
We begin by showing that the true nuisance functions $\pi_{1}$ and $\pi_{2}$ satisfy conditions (1) and (2) with equality, as stated in the following result.
@@ 5.1 Bias Analysis and Robustness Properties
    $E\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]=0,~~~~~~\forall h_{1}\in L^{2}(P_{0,X}).$
@@
    $E\left[\left\{A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]=0,~~~~~~\forall h_{2}\in L^{2}(P_{0,(M,X)}).$
Suppose conditions (1) and (2) hold with equality and $\mathcal{H}_{1}=L^{2}(P_{0,X})$ and $\mathcal{H}_{2}=L^{2}(P_{0,(M,X)})$. Then, it is easy to see that $\hat{\pi}_{1}=\pi_{1}$ and $\hat{\pi}_{2}=\pi_{2}$ with probability one. However, if $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are restricted hypothesis spaces rather than the full $L^{2}$ spaces, conditions (1) and (2) may still hold with equality even though $\hat{\pi}_{1}\neq\pi_{1}$ and $\hat{\pi}_{2}\neq\pi_{2}$. Nevertheless, satisfying conditions (1) and (2) can be sufficient for controlling the bias of the final estimator of the parameter of interest. We formalize this argument by looking at the structure of the bias of $\hat{\psi}^{\text{2S}}$.
@@
The bias of the estimator $\hat{\psi}^{\text{2S}}$ satisfies the following equality:
$E[\hat{\psi}^{\text{2S}}]-\psi_{0}$ $=E\left[\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~+E\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
Note that our estimator $\hat{\psi}^{\text{2S}}$ is based on an influence function, which always leads to a second-order bias. From the structure of the bias of the estimator $\hat{\psi}^{\text{2S}}$, it is evident that if there exists $h_{1}\in\mathcal{H}_{1}$ which closely approximates $\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}$, and there exists $h_{2}\in\mathcal{H}_{2}$ which closely approximates $\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}$, and further conditions (1) and (2) hold, then we can eliminate the bias of the estimator $\hat{\psi}^{\text{2S}}$ even if we do not have a perfect fit of $\pi_{1}$ and $\pi_{2}$. This is in fact the main motivation of our two-stage approach. It substitutes fitting $p(A=1\mid X)$ and $p(M\mid A,X)$, which can be a difficult task, by a much weaker requirement based on balancing.
We, next, analyze the robustness of our proposed method against misspecification of the nuisance functions. Under the following positivity assumption, our approach possesses a multiple-robustness property, as stated in Proposition 6.
@@
For some $\epsilon>0$, we have $p(A=a\mid X=x)>\epsilon$ and $p(M=m\mid A=a,X=x)>\epsilon$, for all $m$, $a$, and $x$.
@@
Under Assumption 2 , $\hat{\psi}^{\text{2S}}$ is asymptotically unbiased if at least one of the following conditions holds:.
  1. (i)
@@
Figure 2 summarizes the result in Proposition 6.
$\hat{\pi}_{1}$$\hat{\pi}_{2}$$\hat{\mu}_{1}$$\hat{\mu}_{2}$Stage 1:Stage 2: